1. I have used the PMC trans beam in my set-up as the required PSL beam.
2. I have superposed the ETMX-Fibre output with the PSL beam on the PSL table.
3. I have used suitable beam splitters and lens to match the power and the sizes of the overlapping beams and have aligned them to the optimum.
4. A lens having f=7.6 cms is used to focus the beam into the PD.
5. Initially, I used the broadband 1611 NewFocus PD to find the IR beat signal by scanning the oven temperature. (using the digital sitemap controls.)
6. I checked the previous elog entries by Suresh and Koji on the green beat signal they had worked on and used their data to get an idea of the temperature range of the oven where I could obtain a beat.
7. I obtained peaks at three different temperatures as had been noted previously and set the temperature so that I am now sitting in the middle stable regime.
8. Then I switched to the 1811 100 MHz PD as it has a larger gain. It has a saturation power of 100 microWatts. The input power at the PD is measured to be 80 microWatts.
9. I was having trouble getting a clean peak due to presence of many harmonics as seen on the spectrum analyser. This happened because there was too much power incident on the PD which led to arising of non-linearity giving rise to harmonics.
10.To reduce the power entering the PD, I put in a ND 1.0 Filter just before the beam enters the PD and obtained a clean signal.
11. I will use the frequency counter tomorrow to check the resonant frequency and try to connect the output to acquire a digital signal.
12. Otherwise I will proceed to build a Mixer Frequency Discriminator.
13. After the feed-back loop is completed, I will proceed to compare the frequency-noises of the green-beat lock and the IR-beat lock.